# On the modeling of irreversibility by relaxator Liouville dynamics

### v1

@@ I Introduction
The irreversibility of macroscopic processes expressed by the basic thermodynamic laws seems to be in contradiction with the reversibility of the microscopic von Neumann equation. The debate of this issue goes back to the early days of statistical mechanics and is still with us (see e.g. Boltzmann1872 -Handbook2011 ). To stick to the von Neumann equation and modify e.g. the definition of entropy and/or refer to the role of special initial conditions that are rare in some probabilistic sense cannot suffice to recover macroscopic irreversibility, because reversible motions cannot relax to a steady state in a manner independent of the initial state. The dynamics should tell us about the relaxation time $\tau_{r}$ to reach stationarity. To extract relaxation as an emerging qualitative feature from the von Neumann equation it needs a treatment under the condition of macroscopic $N$. It is long known that microscopic reversibility has one consequence which must be sacrificed for macroscopic irreversibility: recurrence of initial states after some recurrence time $T_{R}$. Thus, there is one obvious condition for macroscopic irreversibility, already formulated by Smoluchowski in 1916 Smoluchowski1916 : the available time $t_{s}$ to explore the system has to be negligible compared to the recurrence time $T_{R}$. Typically, recurrence times grow double exponential with $N$ and quickly reach super-astronomic time scales for many body systems (for a detailed discussion see Ven2015 ).
@@
As we will argue now, the essential condition for irreversibility in macroscopic systems is the smallness of $t_{s}$ with respect to the time scale $t_{e}$ that is needed to explore the full spectral properties of all degrees of freedom involved. In macroscopic systems $t_{e}$ scales exponentially with $N$ (see below) while $t_{s}$ does not depend on $N$. The condition $t_{s}\ll t_{e}$ guarantees that the Smoluchowski criterion is fulfilled with ease. The spectral exploration time $t_{e}$ is the inverse of the typical spacing of eigenfrequencies $\Omega_{n}$ of $L$ in a certain dynamic regime of frequencies $\Omega$. The condition $t_{s}\ll t_{e}$ means that we can treat the spectrum of $L$ as continuous with a spectral density $\gamma_{L}(\Omega)$, because we cannot resolve it when following the system over reasonable time scales of order $t_{s}$. One immediate consequence is that recurrence times are treated as infinite, because in a continuum of periods there is no finite common least multiplier. A second consequence corresponds to the relaxation of time dependent correlation functions $c_{A}(t)=\left\langle A(t)A\right\rangle$ of time dependent observables $A(t)$ with respect to a stationary state $\rho$, $c_{A}(t)=\sum_{n}P^{[A]}_{n}e^{\imath\Omega_{n}t}$, with $P^{[A]}_{n}>0$. Such superposition with lots of different frequencies leads to almost cancellations and thus irreversible relaxation to a stationary value in a macroscopic system under a limited spectral resolution. With full spectral resolution of all terms in the superposition, one could find almost recurrence of some initial value after $T_{R}$, but on time scales $t_{s}\ll t_{e}\ll T_{r}$ the relaxation to stationarity is the appropriate picture Ven2015 . Indeed, to take imperfect resolution into account one approximates the sum in $c_{A}(t)$ with an integral over the density $\gamma_{L}(\Omega)$, with $\gamma_{L}(\Omega)\cdot P^{[A]}(\Omega)$ of a finite range of order $1/\tau_{A}$. Then one finds the correlations to relax to stationarity on a scale of order $\tau_{A}$. Thus, the condition of $t_{s}\ll t_{e}$ in macroscopic systems is the reason for treating the spectrum of all degrees of freedom as continuous with a spectral function $\gamma_{L}(\Omega)$. With the continuity of $\gamma_{L}(\Omega)$ macroscopic irreversibility shows up in correlation functions. In addition, the continuity and smoothness assumption of the spectral density is necessary in the formulation of equilibrium thermodynamics. The very definition of entropy relies on these conditions (see e.g. LL5 ), $S(E)=\ln(\gamma_{H}(E)\cdot\Delta)$, where $\gamma_{H}(E)$ is the density of energy eigenvalues at energy $E$ and $\Delta$ is an arbitrary fixed energy window sufficiently separated from the lower value $1/t_{e}$ and the upper value $1/t_{s}$. Only then, temperature $T(E)$ can be identified by $1/T(E):=\partial_{E}S(E)$. Energy as a continuous parameter of the macroscopic thermodynamic equilibrium state manifold could not be defined without the continuity condition of the spectral density $\gamma_{H}(E)$. Thermodynamics with Legendre transform relations between thermodynamic potentials derived from partition sums of different equilibrium ensembles relies on the smoothness condition as well and is justified up to relative fluctuations of order $1/\sqrt{N}$ in the thermodynamic potentials (see e.g. Zia2009 ). Since thermodynamic entropy $S(E)$ is an extensive quantity, the condition of continuity $t_{s}\ll t_{e}$ is easily justified as $t_{e}\propto\gamma_{L}(E)\propto e^{S(E)}\propto e^{N}$ which is exponentially larger in $N$ than $t_{s}$. Since the spectral densities $\gamma_{H}(E)$ and $\gamma_{L}(\Omega)$ are related by convolution, $\gamma_{L}(\Omega)=(\gamma_{H}(-E)\ast\gamma_{H}(E))(\Omega)$, qualitative considerations of continuity and scaling with $N$ apply equally to both.
@@
is the resolvent of Liouville $L_{\rm tot}$ with poles at the real valued eigenfrequencies $\Omega_{n}$ symmetrically distributed around $\Omega=0$. When the limit to a continuous spectrum of $L_{\rm tot}$ has been taken, the limit $\epsilon\to 0^{+}$ can usually be performed. For systems with a many body localization (see e.g.Abanin2019 ) phase with respect to a certain basis the global spectrum may be dense, but the typical local density of states vanishes for $N\to\infty$. Thus, typically $t_{e}\to 0$ and no equilibration can take place. Such many body localization phases will not be considered further in the course of this work.
@@
In Section III we derive and discuss the general properties of relaxator Liouville dynamics thereby relating $L(\omega)$ to the Hamiltonian of relevant variables interacting with irrelevant variables. Although there are several approaches to non-Markovian irreversible dynamics (see e.g. BreuEtal -ZhangEtAl2012 ) we believe that the relaxator Liouville approach is most general and covers them in a systematic and transparent way, well suited for phenomenological, axiomatic, and constructive modelings. In Section IV kinetic equations (see e.g. Bonitz2016 ) for one-particle densities are constructed as a special application of Liouville relaxator dynamics where one-particle degrees of freedom are taken as the relevant ones. The corresponding relaxator determines the self-energy in the corresponding non-equilibrium Green’s function NEGF (see e.g. VLeuwen2006 ,Datta ). By incorporating driving fields for currents in the system we derive linear response susceptibilities for relaxator Liouville dynamics that generalize Kubo’s linear response theory Kubo1957 and discuss different sources of irreversibility in Sec. V. By a weak coupling approximation in Sec. VI it is possible to express the relaxator Liouville in terms of isolated properties of relevant and irrelevant variables. We investigate the Markov approximation to compare our general approach with the well known semi-group approach after Lindblad and Gorini ( for overviews see e.g. BreuPet -Lidar2020 ) to irreversible systems in Sec. VII. Finally we give a summary in Sec. VIII.
@@ II Instability of Reversibility
We consider an arbitrary isolated quantum system denoted as total system that allows to distinguish two groups of degrees of freedom, relevant and irrelevant, there numbers being $N_{s}$ and $N_{e}$, respectively. The total Hilbert space is the product of the Hilbert spaces of relevant and irrelevant variables. We want to model the dynamics of the relevant variables up to a time scale $t_{s}$.
It is now a defining condition for the distinction between relevant and irrelevant degrees of freedom that the total system as well as the idealization of an isolated system of irrelevant variables has a spectral resolution time of the same order, denoted as $t_{e}$, which has to be larger than $t_{s}$ by a factor which is exponential in $N_{e}/N_{s}\gg 1$. The total system has a huge number $N=N_{e}+N_{s}\approx N_{e}$ of degrees of freedom most of which are irrelevant. It has become popular to denote such systems of relevant variables in contact with a huge number of irrelevant variables as open system (in the sense of BreuPet ). The relevant variables are denoted simply as system variables and the irrelevant ones as environment. Note however, that the notion of open system here only means a system of variables which is not isolated but interacts with the environment. It is possible, that such open system represents a closed system in the thermodynamic sense, i.e. a system that even in equilibrium exchanges energy with its environment but no matter. Both sets of variables (system and environment) may act within the same volume (e.g. in Brownian motion or in the scattering of electrons with phonons or with lattice imperfections and impurities) and may be even based on the same microscopic particles (e.g. in the relaxation to stationarity of a fluid of particles by self-interaction). It is not necessary that the system exchanges particles with the environment, but it is also not excluded. The belonging of chosen degrees of freedom to system or environment however is held fixed in the modeling. In case the environment surrounds the volume of the system we call it a reservoir (e.g. a small piece of matter subject to radiation). In case that the reservoir is in thermal equilibrium, we call it a heat bath.
@@
It has the following intuitive interpretation: the first term ${L}_{{\cal P}}$ is an isolated system Liouville in the absence of an environment whereas the second term describes virtual processes in the system triggered by the environment, ${L}_{{\cal P}{\cal Q}}G_{\cal Q}(z){L}_{{\cal Q}{\cal P}}$, hopping to $\cal Q$}-space, taking there a lift with isolated $\cal Q$}-propagator and finally hopping back to $\cal P$}-space, see Fig. 1. This second term is no longer Hermitian and breaks reversibility.
@@
where $\rm P$} stands for the Cauchy value on integration. Due to the $\delta$ functional in (15), the Hermiticity of L(z){L}(z) is still valid for almost all values of $\omega$, except for a number of zero measure. If, however, the spectrum $\Omega_{n}$ of complementary dynamics cannot be resolved on the scale of the system’s exploration time $t_{s}\ll t_{e}$ it has to be treated as distributed with a positive, even spectral density $\gamma_{\small{\cal Q}}(\omega)$ and we arrive at
@@
The presence of the relaxator is the central new element of the dynamics for states in a macroscopic system on time scales $t_{s}\ll t_{e}$. The relaxator is given by a resolvent $G_{\cal Q}(z)$ within a continuous spectral resolution and the hopping super operators. It can also be given in terms of the spectral density for the complementary dynamics, $\gamma_{\small{\cal Q}}(\omega)$, corresponding eigenspace projectors $\Pi^{\cal Q}(\omega)$, and the hopping super operators, $L_{{\cal P}{\cal Q}}$, $L_{{\cal Q}{\cal P}}$ as
@@ III Relaxator Liouville dynamics: general properties
Equation (13) allows to estimate the relevant time scale of relaxation indicating irreversibility. The propagator of the complementary dynamics, $\left[z-{L}_{\cal Q}\right]^{-1}\propto t_{e}(\omega)$, scales with the spectral exploration time $t_{e}(\omega)$ of the complementary dynamics at frequency $\omega$. The couplings ${L}_{{\cal P}{\cal Q}}\propto{L}_{{\cal Q}{\cal P}}\propto r_{\rm se}$ scale with the rates $r_{\rm se}$ for transitions between $\cal P$}- and $\cal Q$}-spaces and thus yield the transition time $\tau_{\rm se}$ as the typical inverse transition rate for states to hop between system and environment. Since the relaxator $\Gamma(\omega)$ introduces relaxation we can estimate the relaxation time at frequency $\omega$ by dimensional analysis as
$\tau_{r}(\omega)=a\cdot\tau_{\rm se}\frac{\tau_{\rm se}}{t_{e}(\omega)}\,$, (21)
with some dimensionless number $a$ of ${\cal O}(1)$.
@@
We expect that the ratio of the two sides of (22) is set by the typical time scales associated with ${L}_{{\cal P}{\cal Q}}$ and with $\left[z-{L}_{\cal Q}\right]^{-1}$, respectively. By dimensional analysis we find as a condition for neglecting initial correlations
$\frac{t_{e}(\omega)}{\tau_{\rm se}}\ll 1\,.$ (23)
@@
where the notation implies the Hilbert-Schmidt scalar product. Since the scalar product of the left unit matrix with the initial state amounts to taking the trace, any influence of the initial state is getting lost in a unique stationary state $\rho_{\infty}$. In the non-generic case of degenerate zero modes (48) will change to a sum of projectors on different eigenstate combinations of zero modes. It is obvious that in such case the scalar product of left eigenoperator with the initial state is not simply unity and these scalar products store information about the initial state. Stable degenerate zero modes must be protected against perturbations by design or by some stabilizing natural mechanism. Only with stable degenerate zero modes initial conditions can influence the final stationary states. Such non-generic scenarios are discussed in Albert2018 for systems with semi-group dynamics.
@@
with an environmental correlation function $\gamma_{kk^{\prime}}(\Omega)$ becomes possible if we can rewrite the total energies $E_{\alpha m}$ as additive with respect to system and environment, $E_{m\alpha}\approx E_{\alpha}+\epsilon_{m}$ and treat interaction energies between system and environment as small compared to the resolution scale $1/t_{s}$. This will be undertaken in Sec VI.

### v2

@@ I Introduction
The irreversibility of macroscopic processes expressed by the basic thermodynamic laws seems to be in contradiction with the reversibility of the microscopic von Neumann equation. The debate of this issue goes back to the early days of statistical mechanics and is still with us (see e.g. Boltzmann1872 -Handbook2011 ). To stick to the von Neumann equation and modify e.g. the definition of entropy and/or refer to the role of special initial conditions that are rare in some probabilistic sense cannot suffice to recover macroscopic irreversibility, because reversible motions cannot relax to a steady state in a manner independent of the initial state. The dynamics should tell us about the relaxation time $\tau_{r}$ to reach stationarity. To extract relaxation as an emerging qualitative feature from the von Neumann equation it needs a treatment under the condition of macroscopic $N$. It is long known that microscopic reversibility has one consequence which must be sacrificed for macroscopic irreversibility: recurrence of initial states after some recurrence time $T_{R}$. Thus, there is one obvious condition for macroscopic irreversibility, already formulated by Smoluchowski in 1916 Smoluchowski1916 : the available time $t_{s}$ up to which relevant degrees of freedom of a system are tracked has to be negligible compared to the recurrence time $T_{R}$. To track relevant degrees of freedom in sufficient detail $t_{s}$ should be of the order of the typical inverse level spacing of an isolated system of only relevant variables. It can thus be calculated as the density of frequency levels of an isolated system of only relevant degrees of freedom. $T_{R}$ can be calculated as the typical time after which the return probability of an initial state reaches again a value significantly higher than its average value. For a time signal as superposition of oscillations with different periods, the time $T_{R}$ is the common least multiplier of the periods. A practical way of calculation is given in Ven2015 . Typically, recurrence times grow double exponential with $N$ and quickly reach super-astronomic time scales for many body systems (for a detailed discussion see Ven2015 ).
@@
As we will argue now, the essential condition for irreversibility in macroscopic systems is the smallness of $t_{s}$ with respect to the spectral resolution time $t_{e}$ that is needed to resolve the full spectral properties of all degrees of freedom involved. $t_{e}$ can be calculated as the density of frequency levels of the total system or of an isolated system of only irrelevant degrees of freedom since the total system is assumed to have a huge number of degrees of freedom almost all of which are irrelevant. By this it is the inverse of the typical spacing of eigenfrequencies of the total system in a certain dynamic regime of frequencies. In macroscopic systems $t_{e}$ scales exponentially with $N$ (see below) while $t_{s}$ does not depend on $N$. The condition $t_{s}\ll t_{e}$ guarantees that the Smoluchowski criterion is fulfilled with ease. The condition $t_{s}\ll t_{e}$ means that we can treat the spectrum of $L$ as continuous with a continuous density of frequency levels $\gamma_{L}(\Omega)$, because we cannot resolve discrete levels when following the system over reasonable time scales of order $t_{s}$. One immediate consequence is that recurrence times are treated as infinite, because in a continuum of periods there is no finite common least multiplier. A second consequence corresponds to the relaxation of time dependent correlation functions $c_{A}(t)=\left\langle A(t)A\right\rangle$ of time dependent observables $A(t)$ with respect to a stationary state $\rho$, $c_{A}(t)=\sum_{n}P^{[A]}_{n}e^{\imath\Omega_{n}t}$, with $P^{[A]}_{n}>0$. Such superposition with lots of different frequencies leads to almost cancellations and thus irreversible relaxation to a stationary value in a macroscopic system under a limited spectral resolution. With full spectral resolution of all terms in the superposition, one could find almost recurrence of some initial value after $T_{R}$, but on time scales $t_{s}\ll t_{e}\ll T_{R}$ the relaxation to stationarity is the appropriate picture Ven2015 . Indeed, to take imperfect resolution into account one approximates the sum in $c_{A}(t)$ with an integral over the density $\gamma_{L}(\Omega)$, with $\gamma_{L}(\Omega)\cdot P^{[A]}(\Omega)$ of a finite range of order $1/\tau_{A}$. Then one finds the correlations to relax to stationarity on a scale of order $\tau_{A}$. Thus, the condition of $t_{s}\ll t_{e}$ in macroscopic systems is the reason for treating the spectrum of all degrees of freedom as continuous with a continuous density of frequency levels $\gamma_{L}(\Omega)$. With the continuity of $\gamma_{L}(\Omega)$ macroscopic irreversibility shows up in correlation functions. In addition, the continuity and smoothness assumption of the density of frequency levels is necessary in the formulation of equilibrium thermodynamics. The very definition of entropy relies on these conditions (see e.g. LL5 ), $S(E)=\ln(\gamma_{H}(E)\cdot\Delta)$, where $\gamma_{H}(E)$ is the density of energy levels at energy $E$ and $\Delta$ is an arbitrary fixed energy window sufficiently separated from the lower value $1/t_{e}$ and the upper value $1/t_{s}$. Only then, temperature $T(E)$ can be identified by $1/T(E):=\partial_{E}S(E)$. Energy as a continuous parameter of the macroscopic thermodynamic equilibrium state manifold could not be defined without the continuity condition of the density $\gamma_{H}(E)$. Thermodynamics with Legendre transform relations between thermodynamic potentials derived from partition sums of different equilibrium ensembles relies on the smoothness condition as well and is justified up to relative fluctuations of order $1/\sqrt{N}$ in the thermodynamic potentials (see e.g. Zia2009 ). Since thermodynamic entropy $S(E)$ is an extensive quantity, the condition of continuity $t_{s}\ll t_{e}$ is easily justified as $t_{e}\propto\gamma_{L}(E)\propto e^{S(E)}$ which is exponentially larger in $N$ than $t_{s}$. Since the densities $\gamma_{H}(E)$ and $\gamma_{L}(\Omega)$ are related by convolution, $\gamma_{L}(\Omega)=(\gamma_{H}(-E)\ast\gamma_{H}(E))(\Omega)$, qualitative considerations of continuity and scaling with $N$ apply equally to both.
@@
is the resolvent of Liouville $L_{\rm tot}$ with poles at the real valued eigenfrequencies $\Omega_{n}$ symmetrically distributed around $\Omega=0$. When the limit to a continuous spectrum of $L_{\rm tot}$ has been taken, the limit $\epsilon\to 0^{+}$ can usually be performed. For systems with a many body localization (see e.g.Abanin2019 ) phase with respect to a certain basis the global spectrum may be dense, but the typical local density of energy levels vanishes for $N\to\infty$. Thus, typically $t_{e}\to 0$ and no equilibration can take place. Such many body localization phases will not be considered further in the course of this work.
@@
In Section III we derive and discuss the general properties of relaxator Liouville dynamics thereby relating $L(\omega)$ to the Hamiltonian of relevant variables interacting with irrelevant variables. Although there are several approaches to non-Markovian irreversible dynamics (see e.g. BreuEtal -ZhangEtAl2012 ) we believe that the relaxator Liouville approach is most general and covers them in a systematic and transparent way, well suited for phenomenological, axiomatic, and constructive modelings. In Section IV kinetic equations (see e.g. Bonitz2016 ) for one-particle densities are constructed as a special application of Liouville relaxator dynamics where one-particle degrees of freedom are taken as the relevant ones. The corresponding relaxator determines the self-energy in the corresponding nonequilibrium Green function NEGF (see e.g. VLeuwen2006 ,Datta ). By incorporating driving fields for currents in the system we derive linear response susceptibilities for relaxator Liouville dynamics that generalize Kubo’s linear response theory Kubo1957 and discuss different sources of irreversibility in Sec. V. By a weak coupling approximation in Sec. VI it is possible to express the relaxator Liouville in terms of isolated properties of relevant and irrelevant variables. We investigate the Markov approximation to compare our general approach with the well known semi-group approach after Lindblad and Gorini ( for overviews see e.g. BreuPet -Lidar2020 ) to irreversible systems in Sec. VII. Finally we give a summary in Sec. VIII.
@@ II Instability of Reversibility
We consider an arbitrary isolated quantum system denoted as total system that allows to distinguish two groups of degrees of freedom, relevant and irrelevant, there numbers being $N_{s}$ and $N_{e}$, respectively. The total Hilbert space is the product of the Hilbert spaces of relevant and irrelevant variables. The relevant variables are tracked up to $t_{s}$ which can be identified as the typical inverse level spacing of an isolated system of only relevant variables. It is now a defining condition for the distinction between relevant and irrelevant degrees of freedom that the total system as well as the idealization of an isolated system of irrelevant variables has a spectral resolution time of the same order, which we denoted as $t_{e}$. It has to be larger than $t_{s}$ by a factor which is exponential in $N_{e}/N_{s}\gg 1$. The total system has a huge number $N=N_{e}+N_{s}\approx N_{e}$ of degrees of freedom almost all of which are irrelevant. It has become popular to denote such systems of relevant variables in contact with a huge number of irrelevant variables as open system (in the sense of BreuPet ). The relevant variables are denoted simply as system variables and the irrelevant ones as environment. Note however, that the notion of open system here only means a system of variables which is not isolated but interacts with the environment. It is possible, that such open system represents a closed system in the thermodynamic sense, i.e. a system that even in equilibrium exchanges energy with its environment but no matter. Both sets of variables (system and environment) may act within the same volume (e.g. in Brownian motion or in the scattering of electrons with phonons or with lattice imperfections and impurities) and may be even based on the same microscopic particles (e.g. in the relaxation to stationarity of a fluid of particles by self-interaction). It is not necessary that the system exchanges particles with the environment, but it is also not excluded. The belonging of chosen degrees of freedom to system or environment however is held fixed in the modeling. In case the environment surrounds the volume of the system we call it a reservoir (e.g. a small piece of matter subject to radiation). In case that the reservoir is in thermal equilibrium, we call it a heat bath.
@@
It has the following intuitive interpretation: the first term ${L}_{{\cal P}}$ is an isolated system Liouville in the absence of an environment (it may however contain mean-field influences of the environment, see (65) below) whereas the second term describes virtual processes in the system triggered by the environment, ${L}_{{\cal P}{\cal Q}}G_{\cal Q}(z){L}_{{\cal Q}{\cal P}}$, hopping to $\cal Q$}-space, taking there a lift with isolated $\cal Q$}-propagator and finally hopping back to $\cal P$}-space, see Fig. 1. This second term is no longer Hermitian and breaks reversibility.
@@
where $\rm P$} stands for the Cauchy value on integration. Due to the $\delta$ functional in (15), the Hermiticity of L(z){L}(z) is still valid for almost all values of $\omega$, except for a number of zero measure. If, however, the spectrum $\Omega_{n}$ of complementary dynamics cannot be resolved on the scale of the time $t_{s}\ll t_{e}$ it has to be treated as distributed with a positive, even continuous density of frequency levels $\gamma_{\small{\cal Q}}(\omega)$ and we arrive at
@@
The presence of the relaxator is the central new element of the dynamics for states in a macroscopic system on time scales $t_{s}\ll t_{e}$. The relaxator is given by a resolvent $G_{\cal Q}(z)$ within a continuous spectral resolution and the hopping super operators. It can also be given in terms of the continuous density of frequency levels, $\gamma_{\small{\cal Q}}(\omega)$, corresponding eigenspace projectors $\Pi^{\cal Q}(\omega)$, and the hopping super operators, $L_{{\cal P}{\cal Q}}$, $L_{{\cal Q}{\cal P}}$ as
@@ III Relaxator Liouville dynamics: general properties
Equation (13) allows to estimate the relevant time scale of relaxation indicating irreversibility. The couplings ${L}_{{\cal P}{\cal Q}}\propto{L}_{{\cal Q}{\cal P}}$ can be made dimensionless by multiplication with an appropriate factor, $(\tau_{se})^{-1}L_{PQ}$. This factor defines a time scale $\tau_{se}$ denoted as transition time for transitions between $\cal P$}- and $\cal Q$}-spaces. Since $\left[z-{L}_{\cal Q}\right]^{-1}$ is the Laplace transform of the time evolution operator in $\cal Q$}-space the relaxator is the Fourier transform of a time dependent correlation function of coupling operators entering in $L_{{\cal P}{\cal Q}}$. Denoting the typical relaxation time of time dependent correlation functions of two environmental observables as environmental correlation time $\tau_{e}$, we can estimate the factor ${\tau_{r}(\omega)}^{-1}$ which makes $\Gamma(\omega)$ dimensionless by dimensional analysis as
$\tau_{r}(\omega)=a\cdot\tau_{\rm se}\frac{\tau_{\rm se}}{\tau_{e}(\omega)}\,$, (21)
with some dimensionless number $a$ of ${\cal O}(1)$ and $\tau_{e}(\omega)=\tau_{e}/(1+(\omega\tau_{e})^{2})$. Because $\Gamma(\omega)$ introduces non-positive imaginary contributions to the eigenvalues of the Liouville, states can show relaxation to stationary situations (see also later in this section). The system’s relaxation time $\tau_{r}$ can be defined as the time scale after which the system’s state of relevant variables reaches its stationary value exponentially fast. As an order of magnitude estimation we can thus identify the system’s relaxation time with $\tau_{r}(\omega\to 0)\approx\tau_{\rm se}({\tau_{\rm se}}/{\tau_{e}})$ by (21).
@@
We expect that the ratio of the two sides of (22) is set by the typical time scales associated with ${L}_{{\cal P}{\cal Q}}$ and with $\left[z-{L}_{\cal Q}\right]^{-1}$, respectively. By similar dimensional analysis as before we find as a condition for neglecting initial correlations
$\frac{\tau_{e}(\omega)}{\tau_{\rm se}}\ll 1\,.$ (23)
@@
where the notation implies the Hilbert-Schmidt scalar product. Since the scalar product of the left unit matrix with the initial state amounts to taking the trace, any influence of the initial state is getting lost in a unique stationary state $\rho_{\infty}$. The relaxation of any initial state to a unique stationary state is at the heart of irreversible phenomenology and common experience. It has been studied within different modelings of irreversible systems classical and quantum, e.g. in vKamp , BreuPet , StefAlm and MyöhEtAl .
In the non-generic case of degenerate zero modes (48) will change to a sum of projectors on different eigenstate combinations of zero modes. It is obvious that in such case the scalar product of left eigenoperator with the initial state is not simply unity and these scalar products store information about the initial state. Stable degenerate zero modes must be protected against perturbations by design or by some stabilizing natural mechanism. Only with stable degenerate zero modes initial conditions can influence the final stationary states. Such non-generic scenarios are discussed in Albert2018 for systems with semi-group dynamics. Dependence of the long time behavior on initial states has also been observed within modelings with Luttinger liquids (see e.g. PerfEtAl -BuchEtAl ). Luttinger liquids represent fixed point models of one-dimensional interacting Fermions within unitary dynamics in the renormalization group sense. But they have an infinite number of conserved quantities which are however not stable against the breaking of unitarity by perturbations which are considered as irrelevant in the usual renormalization group sense of unitary dynamics. Thus, the dependence of the long time behavior on initial conditions is due to the degeneracy of the zero mode caused by constants of motion. The degeneracy and initial state dependence gets finally lost by small perturbations (see e.g. MatAnd -BuchEtAl ).
@@
with an environmental correlation function $\gamma_{kk^{\prime}}(\Omega)$ becomes possible if we can rewrite the total energies $E_{\alpha m}$ as additive with respect to system and environment, $E_{m\alpha}\approx E_{\alpha}+\epsilon_{m}$ and treat interaction energies between system and environment as small compared to the resolution scale $1/t_{s}$. This will be undertaken in Sec. VI.